# The Influence of Code Smells in Efferent Neighbors on Class Stability

### v1

@@ 1. Introduction
Changing source code is the heartbeat of software maintenance. For most maintenance tasks, such as fixing bugs, refactoring, or delivering new features, teams aim to keep edits small and localized because smaller changes cost less, disturb less context, and reduce the risk of unintended side effects (Kamei et al., 2012). A system exhibits low stability when parts of it require large edits during maintenance tasks or are frequently modified across multiple tasks (Olbrich et al., 2009; Santana et al., 2024b). Such instability greatly increases the effort and risk of making changes. It is therefore important to understand which code characteristics drive this instability, enabling teams to address these factors, maintain system stability, and deliver changes effectively and efficiently.
One long-suspected factor is the presence of code smells: surface-level characteristics in source code that may indicate underlying design problems (Fowler, 1999). Code smells are widely believed to degrade maintainability (Fowler, 1999; Olbrich et al., 2010; Bán and Ferenc, 2014). Prior studies report that smelly classes tend to be edited more frequently and with greater change size than their non-smelly peers because of their large, complex, and non-cohesive nature (Santana et al., 2024b; Khomh et al., 2009; Palomba et al., 2018b; Olbrich et al., 2009). However, when developers change a class, it is not always because the class itself requires modification. A class may change because the classes it depends on via outgoing dependencies have changed, which can create ripple effects along static dependencies, even if the class itself is not smelly (Salama et al., 2019). We refer to these dependent classes as the class’s efferent neighbors [^1] Here, efferent is used as an adjective meaning “conducting outward” (i.e., outgoing), to emphasize that these neighbors are reached via outgoing dependencies from the class. Definition source: https://www.merriam-webster.com/dictionary/efferent. Figure 1 illustrates the concept: class $C$ has outgoing dependencies on classes $N_{1}$ and $N_{2}$, which we treat as $C$’s efferent neighbors. Changes in $N_{1}$ or $N_{2}$ can propagate to $C$ through these dependencies, creating ripple effects that may require $C$ to change as well. Consequently, when the efferent neighbors are smelly, these ripple effects may occur more frequently, thereby reducing the class’s stability. Focusing only on smells within the class captures the first situation but overlooks dependency-driven changes. To assess whether code smells also influence the stability of a class through this second pathway, we must analyze the smells present in its efferent neighbors and explicitly account for the static dependencies between classes.
@@
In addition, prior work suggests that code smells rarely occur in isolation. Multiple smells frequently appear together within the same class or in classes connected by static dependencies (Palomba et al., 2018a; Yamashita et al., 2015). This phenomenon is referred to as code smell interrelation, and multiple code smells can collectively exacerbate maintainability issues (Abbes et al., 2011; Yamashita and Moonen, 2013; Yamashita et al., 2015). A further consequence of this interrelation is code smell interaction, in which two smells are directly connected through static dependencies. These interactions may further complicate maintainability by affecting modularity, comprehensibility, and modifiability because of the direct structural connection between smelly artifacts (Yamashita and Moonen, 2013; Zhang et al., 2025). However, to the best of our knowledge, few studies have empirically examined how code smell interrelation and interaction influence code quality. Therefore, the goal of this confirmatory study is to assess the association between the presence of code smells in a class’s efferent neighbors and its stability, while accounting for the effects introduced by the presence of multiple code smells, that is, code smell interrelation and code smell interaction.
@@ 2.1. Code Smells
Code smells are surface-level characteristics in source code that can indicate deeper design issues (Fowler, 1999). There are different types of code smells, and they can be classified into class-level smells and method-level smells, as some affect the quality of a single method while others impact an entire class (Lanza and Marinescu, 2007). In this study, we focus on ten types of code smells (Lanza and Marinescu, 2007). The method-level smells are Feature Envy (FE), Brain Method (BM), Dispersed Coupling (DiCo), Intensive Coupling (IC), and Shotgun Surgery (SS). The class-level smells are God Class (GC), Brain Class (BC), Data Class (DC), Refused Bequest (RB), and Tradition Breaker (TB). Table 1 details their definitions (Lanza and Marinescu, 2007).
Table 1. Definitions of Code Smells Considered in This Study (Lanza and Marinescu, 2007)
@@
Although artifacts may exhibit additional code smells beyond the ten we consider, our selection follows established practice in software engineering research. We focus on a subset of smells that (i) offer balanced coverage across both levels, (ii) include smells that are widely studied in software maintenance research, occur frequently in codebases, and have been linked to reduced maintainability (Li and Shatnawi, 2007; Cairo et al., 2018; Tufano et al., 2017; Palomba et al., 2018b; Kaur, 2020; Reis et al., 2020; Zakeri-Nasrabadi et al., 2023), (iii) have clear definitions and widely accepted detection techniques (Lanza and Marinescu, 2007), (iv) are supported by validated detection tools (Fernandes et al., 2016), and (v) align with prior studies, enabling meaningful comparison (Santana et al., 2024b). Our goal is not to exhaustively cover all possible smells, but to analyze a representative subset. Consequently, we believe this choice makes our results more interpretable and practical.
We do not include a few well-known code smells such as Long Method (i.e., a method that is too long (Fowler, 1999)) and Large Class (i.e., a class that is too big (Fowler, 1999)) in this study. Although widely discussed, these smells are usually detected using a single aspect, such as method or class length, which on its own is often an overly sensitive heuristic (e.g., a Large Class that is fully cohesive and has a clear single responsibility, or a Long Method that is not deeply nested, accesses only a limited number of variables, and is easy to understand). In contrast, the smells we analyze are defined based on multiple aspects (e.g., the definitions of Brain Method and God Class include length and complexity, but also consider other aspects such as cohesion and nesting level), using richer structural and coupling-related characteristics that are more directly aligned with our dependency-based theory. Therefore, we focus on the ten well-defined smells to strengthen construct validity by aligning smell definitions with our dependency-based theory, while reducing detection noise from single-aspect, overly sensitive heuristics.
@@ 2.2. The Influence of Code Smells on Code Stability
Software stability is a sub-attribute of maintainability (International Organization for Standardization, 1991). Prior research has examined stability from different perspectives, each focusing on a particular category of software artifact. Salama et al. (Salama et al., 2019) classify these perspectives into four levels, where a level refers to the category of artifact being analysed. The four levels are the code level, the architecture level, the design level, and the requirements level. In this paper, we focus on the code level, as this is where code smells manifest; specifically, we are interested in the stability of class artifacts. At this level, stability is defined as (Salama et al., 2019): (i) intrinsic stability: the ability to remain largely unchanged over time (this captures an intrinsic property of the artifact itself, independent of external pressures such as ripple effects); and (ii) ripple-effect stability: the ability to resist the ripple effect of changes (an example of low ripple-effect stability is when one class changes and the classes depending on it also require modification).
Different metrics exist to measure code stability. Martin (Martin, 1994, 2000) proposes a package instability metric, which is proportional to a package’s efferent coupling, that is, the number of classes outside the package on which the classes inside the package depend. The rationale is that a package whose classes rely on many external classes is more likely to change when any of those dependent classes change. This conceptualization can be extended to the class level using the same rationale, where the efferent coupling of a class corresponds to the number of other classes it depends on (i.e., its efferent neighbors). In software evolution research, code stability is typically measured using two process metrics (Olbrich et al., 2009; Khomh et al., 2009; Santana et al., 2024b): change frequency, which measures how often a software entity is modified within a specific period, and change size, which quantifies the magnitude of each change and is usually expressed in terms of lines of code added or deleted (code churn) during each commit.
Empirical studies generally report that code smells are associated with lower code stability. Several investigations show that smelly classes tend to change more frequently and often with larger modifications, with this effect becoming stronger when multiple smells occur in the same class (Santana et al., 2024b; Khomh et al., 2009; Palomba et al., 2018b; Olbrich et al., 2009). Similar patterns were found in server-side web applications, where smelly files were more frequently updated and exhibited greater change size (Bessghaier et al., 2021). A notable exception is reported by Olbrich et al. (Olbrich et al., 2010), who found that although God and Brain Classes exhibit higher change frequency and more defects in absolute terms, these differences disappear or reverse once class size is controlled for, indicating that size is a confounding variable in observed stability differences.
@@ 2.3. Architectural Smells and Their Evolution
Architectural smells are architecture-level design decisions or structural patterns that negatively impact system quality attributes (Garcia et al., 2009). The key difference between architectural smells and code smells lies in the artifacts in which they manifest. Code smells are defined in terms of code-level (i.e., implementation-level) artifacts, such as methods and classes, whereas architectural smells are defined in terms of architectural-level artifacts, such as components and subsystems (Garcia et al., 2009). For example, the code smell God Class and the architectural smell God Component (Azadi et al., 2019) both describe an artifact that has too many responsibilities. However, God Class refers to an overloaded class, while God Component refers to an overloaded component.
Previous research has studied the evolution of architectural smells. Sas et al. (Sas et al., 2019) tracked how individual smell instances evolve over time by mining three smell types across 14 open-source Java projects. They showed that smell types differ in multiple aspects, such as growth rate and lifetime, offering guidance for smell refactoring and prioritization. Extending this work, Sas et al. (Sas et al., 2022) analyzed how smells evolve in industrial systems. They tracked smells in nine C/C++ projects and complemented mining with interviews with 12 developers. Their findings suggest that smells can persist across releases, overlap within the same components, and be perceived by developers as contributing to maintainability difficulties. Developers also described how smell analysis can provide actionable refactoring insights.
Gnoyke et al. (Gnoyke et al., 2021) studied the evolution of smells and their impact on system degradation by analyzing 14 open-source Java systems. They reported that smells remain mostly stable relative to code size, although some smell types have a greater impact on system degradation. They also identified properties that influence the strength of smells’ impact on system degradation. Recently, Gnoyke et al. (Gnoyke et al., 2024) expanded this work by characterizing typical evolution patterns of smells. They found that some smells tend to accumulate and merge into more complex structures over time, while others remain relatively stable.
@@ 2.4. The Influence of Interrelated and Interacting Code Smells on Maintainability
Code smells are widely recognized and have been empirically shown to undermine maintainability (Fowler, 1999; Olbrich et al., 2010; Bán and Ferenc, 2014). However, they rarely occur in isolation. Instead, multiple code smells often appear together in the same class or in related classes, collectively worsening maintainability. This phenomenon is known as code smell interrelation (Yamashita and Moonen, 2013). Yamashita and Moonen (Yamashita and Moonen, 2013) categorized interrelations as follows: (i) code smell collocation: multiple smells within the same class; and (ii) code smell coupling: smells appearing in different classes that are connected through static dependencies.
Several studies have examined how the collocation of multiple smells within the same class affects maintainability. Abbes et al. (Abbes et al., 2011) found that individual smells showed little measurable impact on comprehension, whereas their collocation significantly increased the time, effort, and errors involved in understanding the code. Later work (Martins et al., 2021; Santana et al., 2024a) showed that classes with multiple collocated smells tend to have higher complexity and lower cohesion, with some combinations, such as a God Class collocated with Feature Envy, being especially detrimental. Developers also reported that collocated smells make code harder to understand and modify (Martins et al., 2021).
Interestingly, prior work suggests that maintenance problems may arise not only from the presence of multiple interrelated smells, but from their interaction through static dependencies (Yamashita, 2012; Yamashita et al., 2015). In an exploratory study, Yamashita and Moonen (Yamashita and Moonen, 2013) observed professional developers performing maintenance tasks and collected their reflections through interviews and think-aloud sessions. Developers indicated that collocated and coupled smells can interact along static dependencies, making code harder to follow and causing ripple effects during changes, which increase effort and defect risk. To the best of our knowledge, this remains the only study that directly links such interactions to maintainability, while later work has mainly identified recurring patterns of collocation and coupling without assessing their impact (Yamashita et al., 2015; Fontana et al., 2015; Walter et al., 2018; Sobrinho and Maia, 2021; Martins et al., 2020).
Recently, Zhang et al. (Zhang et al., 2025) formally defined the term code smell interaction: a code smell instance interacts with another code smell instance if there is a static dependency that directly connects the two instances. They concluded that such interactions are associated with degraded modularity.
@@ 3.1. Dependency
We utilize the following definition of dependency from Stevens et al. (Stevens et al., 1974): “…A dependency is when the functioning of one element A requires the presence of another element B…” We interpret this as a directional connection between two code artifacts, flowing from the source artifact to the target artifact. We capture the requirement between the two through different relations. For example, a method (source artifact type) may create (relation type) a class (target artifact type). We therefore define a dependency type as a triple:
@@
In this paper, we consider five artifact types: class, interface, field, constructor, and method. Following and extending previous work (Yamashita et al., 2015; Zhang et al., 2025), we focus on ten types of relations: call, create, contain, cast, use, throws, return, parameter, extend, and implement.
@@ 3.2.2. RQ2
As Martin (Martin, 1994, 2000) suggests, instability is proportional to efferent coupling. Our theory is that if the efferent neighbors exhibit code smells, the focal class becomes less stable than if the efferent neighbors are clean. This is because prior empirical studies suggest that smelly classes tend to change more frequently (Santana et al., 2024b; Khomh et al., 2009; Palomba et al., 2018b; Olbrich et al., 2009), which in turn may lead the classes depending on them to change more often. Such a situation might even be true for non-smelly classes. Therefore, we test both the population of all classes and the population of non-smelly classes to isolate the effect of smells in the focal class. Hence, our RQ2 and its hypotheses are:

### v2

@@ 1. Introduction
Changing source code is the heartbeat of software maintenance. For most maintenance tasks, such as fixing bugs, refactoring, or delivering new features, teams aim to keep edits small and localized because smaller changes cost less, disturb less context, and reduce the risk of unintended side effects (undef, ). A system exhibits low stability when parts of it require large edits during maintenance tasks or are frequently modified across multiple tasks (undefa, ; undefb, ). Such instability greatly increases the effort and risk of making changes. It is therefore important to understand which code characteristics drive this instability, enabling teams to address these factors, maintain system stability, and deliver changes effectively and efficiently.
One long-suspected factor is the presence of code smells: surface-level characteristics in source code that may indicate underlying design problems (undefc, ). Code smells are widely believed to degrade maintainability (undefc, ; undefd, ; undefe, ). Prior studies report that smelly classes tend to be edited more frequently and with greater change size than their non-smelly peers because of their large, complex, and non-cohesive nature (undefb, ; undeff, ; undefg, ; undefa, ). However, when developers change a class, it is not always because the class itself requires modification. A class may change because the classes it depends on via outgoing dependencies have changed, which can create ripple effects along static dependencies, even if the class itself is not smelly (undefh, ). We refer to these dependent classes as the class’s efferent neighbors [^1] Here, efferent is used as an adjective meaning “conducting outward” (i.e., outgoing), to emphasize that these neighbors are reached via outgoing dependencies from the class. Definition source: https://www.merriam-webster.com/dictionary/efferent. Figure 1 illustrates the concept: class $C$ has outgoing dependencies on classes $N_{1}$ and $N_{2}$, which we treat as $C$’s efferent neighbors. Changes in $N_{1}$ or $N_{2}$ can propagate to $C$ through these dependencies, creating ripple effects that may require $C$ to change as well. Consequently, when the efferent neighbors are smelly, these ripple effects may occur more frequently, thereby reducing the class’s stability. Focusing only on smells within the class captures the first situation but overlooks dependency-driven changes. To assess whether code smells also influence the stability of a class through this second pathway, we must analyze the smells present in its efferent neighbors and explicitly account for the static dependencies between classes.
@@
In addition, prior work suggests that code smells rarely occur in isolation. Multiple smells frequently appear together within the same class or in classes connected by static dependencies (undefi, ; undefj, ). This phenomenon is referred to as code smell interrelation, and multiple code smells can collectively exacerbate maintainability issues (undefk, ; undefl, ; undefj, ). A further consequence of this interrelation is code smell interaction, in which two smells are directly connected through static dependencies. These interactions may further complicate maintainability by affecting modularity, comprehensibility, and modifiability because of the direct structural connection between smelly artifacts (undefl, ; undefm, ). However, to the best of our knowledge, few studies have empirically examined how code smell interrelation and interaction influence code quality. Therefore, the goal of this confirmatory study is to assess the association between the presence of code smells in a class’s efferent neighbors and its stability, while accounting for the effects introduced by the presence of multiple code smells, that is, code smell interrelation and code smell interaction.
@@ 2.1. Code Smells
Code smells are surface-level characteristics in source code that can indicate deeper design issues (undefc, ). There are different types of code smells, and they can be classified into class-level smells and method-level smells, as some affect the quality of a single method while others impact an entire class (undefn, ). In this study, we focus on ten types of code smells (undefn, ). The method-level smells are Feature Envy (FE), Brain Method (BM), Dispersed Coupling (DiCo), Intensive Coupling (IC), and Shotgun Surgery (SS). The class-level smells are God Class (GC), Brain Class (BC), Data Class (DC), Refused Bequest (RB), and Tradition Breaker (TB). Table 1 details their definitions (undefn, ).
Table 1. Definitions of Code Smells Considered in This Study (undefn, )
@@
Although artifacts may exhibit additional code smells beyond the ten we consider, our selection follows established practice in software engineering research. We focus on a subset of smells that (i) offer balanced coverage across both levels, (ii) include smells that are widely studied in software maintenance research, occur frequently in codebases, and have been linked to reduced maintainability (undefo, ; undefp, ; undefq, ; undefg, ; undefr, ; undefs, ; undeft, ), (iii) have clear definitions and widely accepted detection techniques (undefn, ), (iv) are supported by validated detection tools (undefu, ), and (v) align with prior studies, enabling meaningful comparison (undefb, ). Our goal is not to exhaustively cover all possible smells, but to analyze a representative subset. Consequently, we believe this choice makes our results more interpretable and practical.
We do not include a few well-known code smells such as Long Method (i.e., a method that is too long (undefc, )) and Large Class (i.e., a class that is too big (undefc, )) in this study. Although widely discussed, these smells are usually detected using a single aspect, such as method or class length, which on its own is often an overly sensitive heuristic (e.g., a Large Class that is fully cohesive and has a clear single responsibility, or a Long Method that is not deeply nested, accesses only a limited number of variables, and is easy to understand). In contrast, the smells we analyze are defined based on multiple aspects (e.g., the definitions of Brain Method and God Class include length and complexity, but also consider other aspects such as cohesion and nesting level), using richer structural and coupling-related characteristics that are more directly aligned with our dependency-based theory. Therefore, we focus on the ten well-defined smells to strengthen construct validity by aligning smell definitions with our dependency-based theory, while reducing detection noise from single-aspect, overly sensitive heuristics.
@@ 2.2. The Influence of Code Smells on Code Stability
Software stability is a sub-attribute of maintainability (undefv, ). Prior research has examined stability from different perspectives, each focusing on a particular category of software artifact. Salama et al. (undefh, ) classify these perspectives into four levels, where a level refers to the category of artifact being analysed. The four levels are the code level, the architecture level, the design level, and the requirements level. In this paper, we focus on the code level, as this is where code smells manifest; specifically, we are interested in the stability of class artifacts. At this level, stability is defined as (undefh, ): (i) intrinsic stability: the ability to remain largely unchanged over time (this captures an intrinsic property of the artifact itself, independent of external pressures such as ripple effects); and (ii) ripple-effect stability: the ability to resist the ripple effect of changes (an example of low ripple-effect stability is when one class changes and the classes depending on it also require modification).
Different metrics exist to measure code stability. Martin (undefw, ; undefx, ) proposes a package instability metric, which is proportional to a package’s efferent coupling, that is, the number of classes outside the package on which the classes inside the package depend. The rationale is that a package whose classes rely on many external classes is more likely to change when any of those dependent classes change. This conceptualization can be extended to the class level using the same rationale, where the efferent coupling of a class corresponds to the number of other classes it depends on (i.e., its efferent neighbors). In software evolution research, code stability is typically measured using two process metrics (undefa, ; undeff, ; undefb, ): change frequency, which measures how often a software entity is modified within a specific period, and change size, which quantifies the magnitude of each change and is usually expressed in terms of lines of code added or deleted (code churn) during each commit.
Empirical studies generally report that code smells are associated with lower code stability. Several investigations show that smelly classes tend to change more frequently and often with larger modifications, with this effect becoming stronger when multiple smells occur in the same class (undefb, ; undeff, ; undefg, ; undefa, ). Similar patterns were found in server-side web applications, where smelly files were more frequently updated and exhibited greater change size (undefy, ). A notable exception is reported by Olbrich et al. (undefd, ), who found that although God and Brain Classes exhibit higher change frequency and more defects in absolute terms, these differences disappear or reverse once class size is controlled for, indicating that size is a confounding variable in observed stability differences.
@@ 2.3. Architectural Smells and Their Evolution
Architectural smells are architecture-level design decisions or structural patterns that negatively impact system quality attributes (undefz, ). The key difference between architectural smells and code smells lies in the artifacts in which they manifest. Code smells are defined in terms of code-level (i.e., implementation-level) artifacts, such as methods and classes, whereas architectural smells are defined in terms of architectural-level artifacts, such as components and subsystems (undefz, ). For example, the code smell God Class and the architectural smell God Component (undefaa, ) both describe an artifact that has too many responsibilities. However, God Class refers to an overloaded class, while God Component refers to an overloaded component.
Previous research has studied the evolution of architectural smells. Sas et al. (undefab, ) tracked how individual smell instances evolve over time by mining three smell types across 14 open-source Java projects. They showed that smell types differ in multiple aspects, such as growth rate and lifetime, offering guidance for smell refactoring and prioritization. Extending this work, Sas et al. (undefac, ) analyzed how smells evolve in industrial systems. They tracked smells in nine C/C++ projects and complemented mining with interviews with 12 developers. Their findings suggest that smells can persist across releases, overlap within the same components, and be perceived by developers as contributing to maintainability difficulties. Developers also described how smell analysis can provide actionable refactoring insights.
Gnoyke et al. (undefad, ) studied the evolution of smells and their impact on system degradation by analyzing 14 open-source Java systems. They reported that smells remain mostly stable relative to code size, although some smell types have a greater impact on system degradation. They also identified properties that influence the strength of smells’ impact on system degradation. Recently, Gnoyke et al. (undefae, ) expanded this work by characterizing typical evolution patterns of smells. They found that some smells tend to accumulate and merge into more complex structures over time, while others remain relatively stable.
@@ 2.4. The Influence of Interrelated and Interacting Code Smells on Maintainability
Code smells are widely recognized and have been empirically shown to undermine maintainability (undefc, ; undefd, ; undefe, ). However, they rarely occur in isolation. Instead, multiple code smells often appear together in the same class or in related classes, collectively worsening maintainability. This phenomenon is known as code smell interrelation (undefl, ). Yamashita and Moonen (undefl, ) categorized interrelations as follows: (i) code smell collocation: multiple smells within the same class; and (ii) code smell coupling: smells appearing in different classes that are connected through static dependencies.
Several studies have examined how the collocation of multiple smells within the same class affects maintainability. Abbes et al. (undefk, ) found that individual smells showed little measurable impact on comprehension, whereas their collocation significantly increased the time, effort, and errors involved in understanding the code. Later work (undefaf, ; undefag, ) showed that classes with multiple collocated smells tend to have higher complexity and lower cohesion, with some combinations, such as a God Class collocated with Feature Envy, being especially detrimental. Developers also reported that collocated smells make code harder to understand and modify (undefaf, ).
Interestingly, prior work suggests that maintenance problems may arise not only from the presence of multiple interrelated smells, but from their interaction through static dependencies (undefah, ; undefj, ). In an exploratory study, Yamashita and Moonen (undefl, ) observed professional developers performing maintenance tasks and collected their reflections through interviews and think-aloud sessions. Developers indicated that collocated and coupled smells can interact along static dependencies, making code harder to follow and causing ripple effects during changes, which increase effort and defect risk. To the best of our knowledge, this remains the only study that directly links such interactions to maintainability, while later work has mainly identified recurring patterns of collocation and coupling without assessing their impact (undefj, ; undefai, ; undefaj, ; undefak, ; undefal, ).
Recently, Zhang et al. (undefm, ) formally defined the term code smell interaction: a code smell instance interacts with another code smell instance if there is a static dependency that directly connects the two instances. They concluded that such interactions are associated with degraded modularity.
@@ 3.1. Dependency
We utilize the following definition of dependency from Stevens et al. (undefam, ): “…A dependency is when the functioning of one element A requires the presence of another element B…” We interpret this as a directional connection between two code artifacts, flowing from the source artifact to the target artifact. We capture the requirement between the two through different relations. For example, a method (source artifact type) may create (relation type) a class (target artifact type). We therefore define a dependency type as a triple:
@@
In this paper, we consider five artifact types: class, interface, field, constructor, and method. Following and extending previous work (undefj, ; undefm, ), we focus on ten types of relations: call, create, contain, cast, use, throws, return, parameter, extend, and implement.
@@ 3.2.2. RQ2
As Martin (undefw, ; undefx, ) suggests, instability is proportional to efferent coupling. Our theory is that if the efferent neighbors exhibit code smells, the focal class becomes less stable than if the efferent neighbors are clean. This is because prior empirical studies suggest that smelly classes tend to change more frequently (undefb, ; undeff, ; undefg, ; undefa, ), which in turn may lead the classes depending on them to change more often. Such a situation might even be true for non-smelly classes. Therefore, we test both the population of all classes and the population of non-smelly classes to isolate the effect of smells in the focal class. Hence, our RQ2 and its hypotheses are: